Yolanda wants 14L of fruit punch that is 10% juice. At the store she finds punch that is 15% juice and punch that is 8% juice. How much of each should she purchase.
Answer by mananth(16949)

Yolanda wants 14L of fruit punch that is 10% juice. At the store she finds punch that is 15% juice and punch that is 8% juice. How much of each should she purchase

let 15% punch required be x
then 8% punch will be 14-x
15% ------------------ 8% -----------------10%
x-------------------==== 14-x-------------- 14
0.15x+0.08(14-x)=0.1*14
0.15x+1.12 - 0.08x = 1.4
0.07x= 1.4- 1.12
0.07x= 0.28
x=0.28/0.07
x=4
15% punch required will be 4 liters and 8% punch required will be 10 liters